how much % of NaCl is required to dissolve in water to make its freezing point 4degree centigrade?
1
Expert's answer
2012-10-01T10:00:40-0400
dT=k*n
k=1.84
dT=4
n=4/1.84=2.17 mol/L
m=2.17*58.5=126.95
g
W=126.95/(126.95+1000)*100%=11.3%